OpenStudy (anonymous):

[9.01] What is the vertex for the graph of y = 2x2 + 12x + 10? (−3, −8) (−3, −44) (3, 64) (3, 58)

OpenStudy (nurali):

y = 2x^2 + 12x + 10 = 2 (x + 1) (x + 5) so this parabola intersects the x axis at x=-1 and x=-5. The midpoint between them is the x-coordinate of the vertex: [(-1) + (-5)] / 2 = -3 and the y-coordinate is given by the function at that point: y = 2 * (-3)^2 + 12 (-3) + 10 = 18 - 36 + 10 = -8 The vertex is (-3, -8).

OpenStudy (agent0smith):

For an equation y = ax^2 + bx + c, find the x-coordinate of the vertex using\[\large x= \frac{ -b }{2a }\] Then plug that number back into y = 2x^2 + 12x + 10 to find the y-coordinate of the vertex.